# VidEdit: Zero-shot and Spatially Aware Text-driven Video Editing

### v3

@@ 4.1 Experimental setup
Dataset. Following (Bar-Tal et al., 2022; Wu et al., 2022; Qi et al., 2023), we evaluate our approach on videos from DAVIS dataset (Pont-Tuset et al., 2017) resized at a $768\times 432$ resolution. The length of these videos ranges from 20 to 70 frames. To automatically create edit prompts, we use a captioning model (Li et al., 2022) to obtain descriptions of the original video content and we manually design 4 editing prompts for each video.
@@
VidEdit setup. Our experiments are based on latent diffusion models (Rombach et al., 2022). We use a version of stable diffusion trained with HED edge detection (Zhang and Agrawala, 2023) at a $512\times 512$ resolution on LAION-5B dataset (Schuhmann et al., 2022). We choose Mask2former (Cheng et al., 2022) as our instance segmentation network. To edit an atlas, we sample pure Gaussian noise (i.e. $\rho=1$) and denoise it for 50 steps with DDIM sampling and classifier-free guidance (Ho and Salimans, 2022). For a single 70 frames video, it takes $\sim 15$ seconds to edit a $512\times 512$ patch in an atlas and $\sim 1$ minute to reconstruct the video with the edit layer on a NVIDIA TITAN RTX. We set up the HED strength $\lambda$ to 1 by default.
@@
Baselines. We compare our method with two text-to-image frame-wise editing approaches and three text-to-video editing baselines. (1) SDEdit (Meng et al., 2021) is a framewise zero-shot editing approach that corrupts an input frame with noise and denoise it with a target text prompt. (2) ControlNet (Zhang and Agrawala, 2023) performs frame-wise editing with an external condition extracted from the target frame. (3) Text2Live (Bar-Tal et al., 2022) is a Neural Layered Atlas (NLA) based method that trains a generator for each text query to optimize a CLIP-based loss. (4) Tune-a-Video (TAV) (Wu et al., 2022) fine-tunes an inflated version of a pre-trained diffusion model on a video to produce similar content. (5) Pix2Video (Ceylan et al., 2023) uses a structure-guided image diffusion model to perform text-guided edits on a key frame and propagate the changes to the future frames via self-attention feature injection.
Metrics. A video edit is expected (1) to faithfully render a target text query, (2) to preserve out-of-interest regions unaltered and (3) to be temporally consistent. First, in order to evaluate the faithfulness of an edit, we report semantic metrics computed with ViT-L/14 CLIP (Radford et al., 2021). Prompt Consistency ($\mathcal{C}_{\text{\scriptsize{Prompt}}}$) measures the CLIP similarity between a text query and each video frame. Frame Accuracy ($\mathcal{A}_{\text{\scriptsize{Frame}}}$) corresponds to the percentage of edited frames that have a higher CLIP similarity with the target prompt than with the source caption. Directional Similarity ($\mathcal{S}_{\text{\scriptsize{Dir}}}$) quantifies how closely the alterations made to an original image align with the changes between a source caption and a target caption. Second, we evaluate original content preservation with image similarity metrics: LPIPS (Zhang et al., 2018), HaarPSI (Reisenhofer et al., 2018) and PSNR. While LPIPS evaluates the perceptual similarity between two images in a deep feature space, HaarPSI performs a Haar wavelet decomposition to assess local similarities. On the contrary, PSNR measures the distance with an original image in the pixel space. Finally, we assess the temporal consistency with Frame Consistency ($\mathcal{C}_{\text{\scriptsize{Frame}}}$) that measures the CLIP similarity between consecutive video frames. We also provide an Aggregate score for semantic and similarity metrics that reflect the overall performance of each model on these aspects. More details on these metrics are provided in Supplementary A.
@@ 4.2 State-of-the-art comparison
Quantitative results. Table 1 gathers the overall comparison with respect to the chosen baselines [^1] Results in bold correspond to the best methods based on a paired $t$-test (risk 5%).. We observe that VidEdit outperforms other approaches in terms of both semantic and similarity metrics. Moreover, it demonstrates better temporal consistency compared to other methods. Additionally, when comparing the processing time of different baselines, we found that VidEdit has a significant advantage, with a $\sim$ 30-fold speed-up factor over Text2Live. We further illustrate the lightweight aspect of our method in Supplementary E. Regarding semantic metrics, as indicated by our best directional similarity score, VidEdit performs highly consistent edits with respect to the change between the target text query and the source caption. Even though our method is close to Text2Live in terms of frame accuracy and prompt consistency, the latter explicitly optimizes a generator on a CLIP-based loss, making the aforementioned metrics not reliable to assess its generalization performance and editing quality, as will be shown in the qualitative results. When it comes to image preservation evaluated with our similarity metrics, VidEdit outperforms all baselines in LPIPS and HaarPSI, and is similar to Text2Live on PSNR. This shows the capacity of our approach to optimally preserve the visual content of the source video while generating faithful edits to the target queries. Finally, VidEdit outperforms all methods in $\mathcal{C}_{\text{\scriptsize{Frame}}}$, and shows that the fine spatial control of our approach also translates in an improved temporal consistency.
@@
Semantic Similarity Temporal Time Method $\mathcal{C}_{\text{\scriptsize{Prompt}}}$($\uparrow$) $\mathcal{A}_{\text{\scriptsize{Frame}}}$($\uparrow$) $\mathcal{S}_{\text{\scriptsize{Dir}}}$($\uparrow$) Agg. Score ($\downarrow$) LPIPS ($\downarrow$) HaarPSI ($\uparrow$) PSNR ($\uparrow$) Agg. Score ($\downarrow$) $\mathcal{C}_{\text{\scriptsize{Frame}}}$($\uparrow$) min VidEdit (ours) 28.1 (±3.0) 91.5 (±11.1) 21.7 (±8.4) 3.06 0.077 (±0.054) 0.730 (±0.109) 22.6 (±3.6) 3.01 97.4 (±1.4) 118 Text2Live (Bar-Tal et al., 2022) 28.7 (±2.8) 94.1 (±14.6) 20.4 (±6.0) 3.07 0.155 (±0.035) 0.710 (±0.088) 22.8 (±2.9) 4.04 97.0 (±1.4) 3780 ControlNet (Zhang and Agrawala, 2023) 28.0 (±2.6) 84.8 (±24.0) 18.7 (±6.6) 3.30 0.647 (±0.061) 0.312 (±0.036) 10.8 (±1.5) 12.85 86.2 (±3.6) 561 SDEdit (Meng et al., 2021) 26.1 (±2.9) 65.7 (±31.8) 14.2 (±7.4) 3.96 0.490 (±0.051) 0.377 (±0.034) 17.9 (±1.6) 9.57 83.9 (±5.2) 303 TAV (Wu et al., 2022) 27.5 (±3.1) 73.4 (±36.3) 13.7 (±9.0) 3.92 0.584 (±0.079) 0.274 (±0.060) 13.0 (±2.1) 12.00 96.4 (±1.6) 274 Pix2Video (Ceylan et al., 2023) 29.0 (±3.0) 82.9 (±30.0) 16.2 (±9.2) 3.47 0.540 (±0.079) 0.326 (±0.069) 13.8 (±2.1) 10.90 94.4 (±2.1) 252
Figure 4: Masked LPIPS vs Local Object Accuracy
@@
To further analyze the fine-grained editing capacity of our method while preserving the original video content, we display for all baselines in Fig. 4, their local $\mathcal{A}_{\text{\scriptsize{Frame}}}$ score computed within a ground truth mask compared to an outer LPIPS metric (denoted O-LPIPS) computed on the invert of the mask [^2] The size of each dot is proportional to the local frame accuracy’s standard deviation.. We see that VidEdit reaches a very good local frame accuracy, even outperforming Text2Live. Morevover, VidEdit shows a huge improvement on the O-LPIPS metric compared to the baselines, including Text2Live (3 vs 8), showing clearly a better preservation of out-of-interest regions.
@@
Qualitative results. We show in Fig. 5 a visual comparison against the baselines to qualitatively assess the improvement brought out by our method. We can see that VidEdit performs fine-grained editing while perfectly preserving out-of-interest regions. In comparison to other baselines, the edits generated are more visually appealing and realistic. For example, VidEdit obtains a frame accuracy ($\mathcal{A}_{\text{\scriptsize{Frame}}}$) and prompt consistency ($\mathcal{C}_{\text{\scriptsize{Prompt}}}$) scores of 26.5 and 30 respectively compared to Text2Live which reaches 26 and 35 respectively. However, we can see that Text2Live’s scores do not automatically translate into high-quality edits it often struggles to render detailed textures precisely localized on targeted regions. For example, ice creams are poorly rendered and some untargeted areas are being altered. Regarding Tune-a-Video and Pix2Video baselines, the methods are unable to generate a faithful edit at the exact location and completely degrade the original content. Despite relatively high frame consistency scores in this video (96% for Tune-a-Video and 89% for Pix2Video vs 97.5% for VidEdit), noticeable flickering artifacts undermine the video content. On the other hand, naive frame-wise application of image-to-image translation methods also leads to temporally inconsistent results. For example, SDEdit is unable to both generate a faithful edit and to preserve the original content as it inherently faces a trade-off between the two. Other visual comparisons are shown in Supplementary C.
@@ 4.3 Model Analysis
Table 2: Ablation study. The mask and the HED map help to generate meaningful edits once mapped back in the original frame space.
@@
In the case where no conditional control is passed on to the model, we observe a substantial drop in semantic metrics as the model generates edits at random locations in the atlas whose shapes don’t match the structure of the target object. The introduction of edge conditioning without spatial awareness is quite similar to the previous case with the difference that the model tries to locally match the control information. This results in slightly better semantic results and similarity metrics than with no edge control. Finally, blending an edit with mask control without taking structure conditioning into account generates an edit at the right location but that is semantically incoherent once mapped back to the original images. Yet, this scenario achieves a decent prompt consistency as the objects still correspond to the target text query. We provide a visual illustration of this ablation study in Supplementary B.
@@
On the left panel, we can see that for $\lambda$ values lower than 0.4, the edge conditioning is not strong enough to guide the edits toward a plausible output on the video frames. This phenomenon is illustrated in Supplementary B. On the contrary, for strength values larger than 1.2, the conditioning weighs too much on the model and hinders its ability to generate faithful edits. As expected, we notice that the local LPIPS decreases as the edge conditioning increases. While the decreasing rate is substantial between 0 and 1, the marginal gain diminishes for larger values. Overall, setting the HED strength between 0.8 and 1.2 robustly enables to both perform faithful edits and preserve the original content. On the right panel, we see that both local CLIP score and LPIPS increase with the noising ratio. Indeed, for a null $\rho$ value, the region is reconstructed from the atlas, nearly identically to the original, and is then rewarded a low LPIPS. However, as no modification has been performed, the patch does not match the target text query and gets a lower CLIP score. As the noising ratio increases, the region deviates more from the input but also better matches the target edit. Note that for a $\rho$ value of 100%, the local LPIPS is constrained below 19, which still indicates a low disparity with the original image.

### v4

@@ 4.1 Experimental setup
Dataset. Following Bar-Tal et al. (2022); Wu et al. (2022); Qi et al. (2023), we evaluate our approach on videos from DAVIS dataset (Pont-Tuset et al., 2017) resized at a $768\times 432$ resolution. The length of these videos ranges from 20 to 70 frames. To automatically create edit prompts, we use a captioning model (Li et al., 2022) to obtain descriptions of the original video content and we manually design 4 editing prompts for each video.
@@
VidEdit setup. To control the semantic layout of our generated edits, we utilize a ControlNet variant of Stable Diffusion (Zhang & Agrawala, 2023). This model has learned to detect and to integrate HED edges as conditional information to a diffusion model via training a copy of its layers while also maintaining locked the pre-trained parameters separately. The trainable and locked copies of the parameters are connected at each block of the UNet decoder via “zero convolution” layers that are also optimized. We refer to the original paper for additional information. The original version of Stable Diffusion is trained at a $512\times 512$ resolution on LAION-5B dataset (Schuhmann et al., 2022). We choose Mask2former (Cheng et al., 2022) as our instance segmentation network. To edit an atlas, we sample pure Gaussian noise (i.e. $\rho=1$) and denoise it for 50 steps with DDIM sampling and classifier-free guidance (Ho & Salimans, 2022). For a single 70 frames video, it takes $\sim 15$ seconds to edit a $512\times 512$ patch in an atlas and $\sim 1$ minute to reconstruct the video with the edit layer on a NVIDIA TITAN RTX, a graphic card accessible to the general public. We set up the HED strength $\lambda$ to 1 by default.
@@
Baselines. We compare our method with two text-to-image frame-wise editing approaches and three text-to-video editing baselines. (1) SDEdit (Meng et al., 2021) is a framewise zero-shot editing approach that corrupts an input frame with noise and denoise it with a target text prompt. (2) ControlNet (Zhang & Agrawala, 2023) performs frame-wise editing with an external condition extracted from the target frame. (3) Text2Live (Bar-Tal et al., 2022) is a Neural Layered Atlas (NLA) based method that trains a generator for each text query to optimize a CLIP-based loss. (4) Tune-a-Video (TAV) (Wu et al., 2022) fine-tunes an inflated version of a pre-trained diffusion model on a video to produce similar content. (5) Pix2Video (Ceylan et al., 2023) uses a structure-guided image diffusion model to perform text-guided edits on a key frame and propagate the changes to the future frames via self-attention feature injection.
Metrics. A video edit is expected to (1) be temporally consistent (Temporal), (2) faithfully render a target text query (Semantics), (3) preserve out-of-interest regions unaltered (Similarity). To evaluate CLIP based metrics, we used CLIP ViT-L/14 (Radford et al., 2021).
  1. $\bullet$
    Temporal
    1. $\cdot$
      Frame Consistency ($\mathcal{C}_{\text{\scriptsize{Frame}}}$). Measures the CLIP similarity between the image embeddings of consecutive video frames. Formally, it writes:
      $\mathcal{C}_{\text{\scriptsize{Frame}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}-1}CLIPScore(\bar{I}_{j}^{k},\bar{I}_{j+1}^{k})$ (5)
      where $\bar{I}_{j}^{k}$ is the $j$-th frame of edited video $k$, $F_{k}$ the number of frames in video $k$ and $N$ the number of edited videos.
    2. $\cdot$
      Warping Error ($\mathcal{E}_{\text{\scriptsize{Warp}}}$). Measures the temporal stability of edited videos based on the flow warping error between two frames:
      $\mathcal{E}_{\text{\scriptsize{Warp}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}-1}Warp(\bar{I}_{j}^{k},\bar{I}_{j+1}^{k})$ (6)
      with $Warp(\bar{I}_{j}^{k},\bar{I}_{j+1}^{k})$, the warping error between consecutive frames of an edited video defined as in Lai et al. (2018)
  2. $\bullet$
    Semantics
    1. $\cdot$
      Prompt Consistency ($\mathcal{C}_{\text{\scriptsize{Prompt}}}$). Measures the average CLIP similarity between a target text query and each video frame. For a unique pair (image; caption) the CLIP similarity writes: $CLIPScore(I,C)=\max(100\times cos(E_{I},E_{C}),0)$ with $E_{I}$ the visual CLIP embedding for an image $I$, and $E_{C}$ the textual CLIP embedding for a caption $C$. $\mathcal{C}_{\text{\scriptsize{Prompt}}}$ then writes :
      $\mathcal{C}_{\text{\scriptsize{Prompt}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}}CLIPScore(\bar{I}_{j}^{k},\bar{C}^{k})$ (7)
      where $\bar{I}_{j}^{k}$ is the $j$-th frame of edited video $k$, $\bar{C}^{k}$ the edited caption of video $k$, $F_{k}$ the number of frames in video $k$ and $N$ the number of edited videos.
    1. $\cdot$
      Frame Accuracy ($\mathcal{A}_{\text{\scriptsize{Frame}}}$). Corresponds to the average percentage of edited frames that have a higher CLIP similarity with the target text query than with their source caption. Formally, $\mathcal{A}_{\text{\scriptsize{Frame}}}$ writes:
      $\mathcal{A}_{\text{\scriptsize{Frame}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}}\mathds{1}\{CLIPScore(\bar{I}_{j}^{k},\bar{C}^{k})>CLIPScore(\bar{I}_{j}^{k},C^{k})\}\times 100$ (8)
    2. $\cdot$
      Directional Similarity ($\mathcal{S}_{\text{\scriptsize{Dir}}}$). Quantifies how closely the alterations made to an original image align with the changes between a source caption and a target caption. For the $j$-th frame of video $k$ the similarity score writes: $SIMScore(\bar{I}_{j}^{k},I_{j}^{k},\bar{C}^{k},C^{k})=100*cos(E_{\bar{I}_{j}^{k}}-E_{I_{j}^{k}},E_{\bar{C}_{j}^{k}}-E_{C_{j}^{k}})$. $\mathcal{S}_{\text{\scriptsize{Dir}}}$ then writes:
      $\mathcal{S}_{\text{\scriptsize{Dir}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{j=1}^{F_{k}}SIMScore\left(\bar{I}_{j}^{k},I_{j}^{k},\bar{C}^{k},C^{k}\right)$ (9)
  3. $\bullet$
    Similarity
    Regarding content preservation, we have chosen three metrics that operate on different feature spaces in order to capture a rich description of perceptual similarities.
    1. $\cdot$
      LPIPS (Zhang et al., 2018) operates on the deep feature space of a VGG network and has been shown to match human perception well.
    2. $\cdot$
      HaarPSI (Reisenhofer et al., 2018) While LPIPS evaluates the perceptual similarity between two images in a deep feature space, HaarPSI performs a Haar wavelet decomposition to assess local similarities.
    3. $\cdot$
      PSNR measures the distance with an original image in the pixel space.
    These metrics are extensively described in the literature and we refer to it for further details.
  4. $\bullet$
@@ 4.2 State-of-the-art comparison
Quantitative results. Tab. 1 gathers the overall comparison with respect to the chosen baselines [^1] Results in bold correspond to the best methods based on a paired $t$-test (risk 5%).. VidEdit outperforms other approaches in terms of both semantic and similarity metrics. Moreover, it exhibits a temporal consistency comparable to Text2Live while largely surpassing alternative approaches. Regarding semantic metrics, as indicated by our best directional similarity score, VidEdit performs highly consistent edits with respect to the change between the target text query and the source caption.
@@
Semantic Similarity Temporal Method $\mathcal{C}_{\text{\scriptsize{Prompt}}}$ ($\uparrow$) $\mathcal{A}_{\text{\scriptsize{Frame}}}$ ($\uparrow$) $\mathcal{S}_{\text{\scriptsize{Dir}}}$ ($\uparrow$) Agg. Score ($\downarrow$) LPIPS ($\downarrow$) HaarPSI ($\uparrow$) PSNR ($\uparrow$) Agg. Score ($\downarrow$) $\mathcal{C}_{\text{\scriptsize{Frame}}}$ ($\uparrow$) $\mathcal{E}_{\text{\scriptsize{Warp}}}$ (×$\times 10^{-3}$) ($\downarrow$) Agg. Score ($\downarrow$) VidEdit (ours) 28.1 (±3.0) 91.5 (±11.1) 21.7 (±8.4) 3.06 0.077 (±0.054) 0.730 (±0.109) 22.6 (±3.6) 3.01 97.4 (±1.4) 5.2 (±9.3) 2.33 Text2Live 28.7 (±2.8) 94.1 (±14.6) 20.4 (±6.0) 3.07 0.155 (±0.035) 0.710 (±0.088) 22.8 (±2.9) 4.04 97.0 (±1.4) 3.9 (±4.9) 2.00 ControlNet 28.0 (±2.6) 84.8 (±24.0) 18.7 (±6.6) 3.30 0.647 (±0.061) 0.312 (±0.036) 10.8 (±1.5) 12.85 86.2 (±3.6) 177.9 (±64.3) 46.74 SDEdit 26.1 (±2.9) 65.7 (±31.8) 14.2 (±7.4) 3.96 0.490 (±0.051) 0.377 (±0.034) 17.9 (±1.6) 9.57 83.9 (±5.2) 71.2 (±30.1) 19.42 TAV 27.5 (±3.1) 73.4 (±36.3) 13.7 (±9.0) 3.92 0.584 (±0.079) 0.274 (±0.060) 13.0 (±2.1) 12.00 96.4 (±1.6) 47.8 (±27.7) 13.27 Pix2Video 29.0 (±3.0) 82.9 (±30.0) 16.2 (±9.2) 3.47 0.540 (±0.079) 0.326 (±0.069) 13.8 (±2.1) 10.90 94.4 (±2.1) 160.9 (±98.5) 42.29
Even though our method is close to Text2Live in terms of frame accuracy and prompt consistency, the latter explicitly optimizes a generator on a CLIP-based loss, making the aforementioned metrics not reliable to assess its generalization performance and editing quality, as will be shown in the qualitative results. When it comes to image preservation evaluated with our similarity metrics, VidEdit outperforms all baselines in LPIPS and HaarPSI, and is similar to Text2Live on PSNR. This shows the capacity of our approach to optimally preserve the visual content of the source video while generating faithful edits to the target queries. Finally, VidEdit outperforms all methods in $\mathcal{C}_{\text{\scriptsize{Frame}}}$, and shows that the fine spatial control of our approach also translates in an improved temporal consistency.
@@
To further analyze the fine-grained editing capacity of our method while preserving the original video content, we display for all baselines in Fig. 4, their local $\mathcal{A}_{\text{\scriptsize{Frame}}}$ score computed within a ground truth mask compared to an outer LPIPS metric (denoted O-LPIPS) computed on the invert of the mask. We see that VidEdit reaches a very good local frame accuracy, even outperforming Text2Live. Morevover, VidEdit shows a huge improvement on the O-LPIPS metric compared to the baselines, including Text2Live (3 vs 8), showing clearly a better preservation of out-of-interest regions.
Figure 4: Masked LPIPS vs Local Object Accuracy. The size of each dot is proportional to the standard deviation of the local object accuracy.
Additionally, when comparing the processing time of different baselines, we found that VidEdit has a significant advantage, with a $\sim$ 30-fold speed-up factor over Text2Live. Focusing on the interative part of editing [^2] Steps as DDIM inversion or LNA construction being considered as pre-processing steps. in which users are interested in [^3] Editing the atlas and reconstructing the video for atlas based methods. Simply inferring the model for other baselines., Figure 5 underlines the lightweight aspect of our method. The panel on the left shows that VidEdit can perform a large number of edits on a 70 frame long video in significantly less time than other appoaches. As an illustration, VidEdit demonstrates approximately 30 times faster editing capabilities compared to Text2Live, which is the second leading baseline in terms of editing capacities. On the other hand, as depicted in the right panel, the use of VidEdit becomes increasingly time-efficient compared to other baselines, as the number of frames to edit growths.
@@
Qualitative results. We show in Fig. 6 a visual comparison against the baselines to qualitatively assess the improvement brought out by our method.We can see that VidEdit performs fine-grained editing while perfectly preserving out-of-interest regions. In comparison to other baselines, the edits generated are more visually appealing and realistic. For example, VidEdit obtains a frame accuracy ($\mathcal{A}_{\text{\scriptsize{Frame}}}$) and prompt consistency ($\mathcal{C}_{\text{\scriptsize{Prompt}}}$) scores of 26.5 and 30 respectively compared to Text2Live which reaches 26 and 35 respectively. However, we can see that Text2Live’s scores do not automatically translate into high-quality edits it often struggles to render detailed textures precisely localized on targeted regions. For example, ice creams are poorly rendered and some untargeted areas are being altered. Regarding Tune-a-Video and Pix2Video baselines, the methods are unable to generate a faithful edit at the exact location and completely degrade the original content. Despite relatively high frame consistency scores in this video (96% for Tune-a-Video and 89% for Pix2Video vs 97.5% for VidEdit), noticeable flickering artifacts undermine the video content. On the other hand, naive frame-wise application of image-to-image translation methods also leads to temporally inconsistent results. For example, SDEdit is unable to both generate a faithful edit and to preserve the original content as it inherently faces a trade-off between the two. Other visual comparisons are shown in Appendix B.
@@ 4.3 Model Analysis
Table 2: Ablation study. Mask and HED map help to generate meaningful edits in the original frame space.
@@
In the case where no conditional control is passed on to the model, we observe a substantial drop in semantic metrics as the model generates edits at random locations in the atlas whose shapes don’t match the structure of the target object. The introduction of edge conditioning without spatial awareness is quite similar to the previous case with the difference that the model tries to locally match the control information. This results in slightly better semantic results and similarity metrics than with no edge control. Finally, blending an edit with mask control without taking structure conditioning into account generates an edit at the right location but that is semantically incoherent once mapped back to the original images. Yet, this scenario achieves a decent prompt consistency as the objects still correspond to the target text query. We provide a visual illustration of this ablation study in Appendix A.
@@
On the left panel, we can see that for $\lambda$ values lower than 0.4, the edge conditioning is not strong enough to guide the edits toward a plausible output on the video frames. This phenomenon is illustrated in Appendix A. On the contrary, for strength values larger than 1.2, the conditioning weighs too much on the model and hinders its ability to generate faithful edits. As expected, we notice that the local LPIPS decreases as the edge conditioning increases. While the decreasing rate is substantial between 0 and 1, the marginal gain diminishes for larger values.
@@
Overall, setting the HED strength between 0.8 and 1.2 robustly enables to both perform faithful edits and preserve the original content. On the right panel, we see that both local CLIP score and LPIPS  increase with the noising ratio. Indeed, for a null $\rho$ value, the region is reconstructed from the atlas, nearly identically to the original, and is then rewarded a low LPIPS. However, as no modification has been performed, the patch does not match the target text query and gets a lower CLIP score. As the noising ratio increases, the region deviates more from the input but also better matches the target edit. Note that for a $\rho$ value of 100%, the local LPIPS is constrained below 19, which still indicates a low disparity with the original image.